A point charge has a charge of 8.00 x 10-11 C. At what distance from the point charge is the electrical potential (a) 12.0 V? (b) 24.0 V?




1
Expert's answer
2022-02-28T10:05:31-0500

Answer.

Distance is given

"r=\frac{kQ}{V}"

a) the distance


"r=\frac{9*10^9*8*10^{-11}}{12}\\=6cm"

b) the potential

"r=\frac{9*10^9*8*10^{-11}}{24}\\=3cm"